slipknot3066:

Hemophilia is a gender-linked blood disorder in humans. The gene responsible for it is gene-linked recessive, meaning it is present on the X chromosome. When the gene is present in males (Xh), it always expresses itself, meaning that the male's blood won't clot properly. When present in females in a heterozygous condition (XhXH), the gene doesn’t express itself; but the female becomes a carrier, meaning that her offspring could have hemophilia.. When present in a homozygous recessive condition in females (XhXh), it expresses itself. One of every 10,000 persons is affected by this disease. What could be the allele frequency for this disease in the entire population?

slipknot3066:

@Ultrilliam @AdiGirl @Aeon

AdiGirl:

wut

AdiGirl:

is there any answer choices

slipknot3066:

nope

AdiGirl:

im sorry but idk this question I haven't studied this yet sorry

Vocaloid:

@BRAINIAC do you know anything about genetics/biology? I was thinking to use Hardy-Weinberg here but not sure if it applies to x-linked traits

BRAINIAC:

The allele frequency of an x-linked recessive disorder is the same as the frequency with which it occurs in men. In real life, the 1 in 10,000 figure is an approximation so we can't meaningfully use it to compute an exact allele frequency, and because the frequency in women is almost negligible, we can say the frequency in men is roughly 1 in 5000 and that that's the approximate allele frequency.

BRAINIAC:

I think u can apply the Hardy-Weinberg eqn here since the question stated that one of every 10,000 persons is affected by this disease. If 1 of 10,000 persons is affected, q2 = 1/10000 = 0.0001 The frequency of hemophilia allele is q = 0.01 p + q = 1, so p = 1 – 0.01 = 0.99 The frequencies of alleles for hemophilia are p = 0.99, and q = 0.01, where p is the negative allele for hemophilia, and q is the positive allele for hemophilia